(16 points) If 1.000 g of chromium metal produces 1.461 g of chromium oxide, what is the empirical formula of the product?
Answer #1

1.00 g of chromium metal produces 1.461 g of chromium oxide

weight of chromium in the sample = 1.00 g

mole of chromium = 1 g /51.9961 g/mol = 0.0192 mol

weight of O in the sample = 1.461-1.00 = 0.461 g

mole of O = 0.461 g/15.999 g/mol = 0.0288 mol

so the mole ratio of chromium and oxygen = 0.0288/0.0192 = 1.5

so in the oxide for 1 mol of chromium 1.5 moles of oxygen will be present

so the empirical formula of the oxide is Cr2O3